# A systematic search for AGN obscuration variability in the Chandra archive

### v1

@@ Abstract
The nature of the obscuring material in active galactic nuclei (AGN) can be studied by measuring changes in the line-of-sight column density, , over time. This can be accomplished by monitoring AGN over long periods of time and at all timescales. However, this can only be done for a few selected objects as it is resource intensive. Therefore, the best option currently is to focus on population statistics based on the available archival data. In this work, we estimate a lower limit on the fraction of sources in the local universe that display spectral variability among observations ( %), indicative of variability. Interestingly, we also find that the variable fraction is similar for both Seyfert 1 ( %) and Seyfert 2 ( %) galaxies, and discuss why we might find a slightly higher . We also present a sample of 43 Seyfert 1 and Seyfert 2 galaxies with multiple Chandra observations whose properties make them promising -variable targets. We discuss the accuracy of our method, and search for potential dependencies on the timescale between variable and non-variable observation pairs within a given source. In agreement with previous studies, we find evidence that more variability occurs on longer timescales than on shorter timescales.
@@ 1 Introduction
Active galactic nuclei (AGN) shine brightly at all wavelengths due to the accretion of matter onto a supermassive black hole (SMBH). The unified model suggests the presence of a dusty, torus-shaped structure that obscures certain lines of sight to the central region, resulting in different AGN classifications depending on the orientation of the system with respect to the observer (e.g. Antonucci, 1993; Urry & Padovani, 1995).
@@
The nature of this obscuring material remains uncertain. Infrared observations, corroborated by theoretical considerations, suggest an inhomogeneous or clumpy medium (e.g. Nenkova et al., 2002, 2008; Gandhi et al., 2009; Stalevski et al., 2012; Alonso-Herrero et al., 2016; García-Bernete et al., 2016). Many ideas have been proposed to explain the origin of the inhomogeneities and maintenance of the geometrical thickness including collisional heating (e.g. Krolik & Begelman, 1988), IR radiation (e.g. Pier & Krolik, 1992; Krolik, 2007), or heating by supernovae in starbursts (e.g. Fabian et al., 1998; Wada & Norman, 2002; Wada et al., 2009). Winds driven by hydromagnetic forces (e.g. Konigl & Kartje, 1994; Elitzur & Shlosman, 2006) or radiation pressure and feedback effects (e.g. Wada, 2012; Schartmann et al., 2014) have also been suggested. All of these mechanisms involve a very complex and dynamic environment, which is also responsible for the obscuration.
@@
X-ray observations further suggest inhomogeneity and dynamism through variability in the line-of-sight hydrogen column density, (e.g. Risaliti et al., 2002). Variability in has been observed on timescales ranging from less than a day (e.g. Elvis et al., 2004; Risaliti et al., 2005; Puccetti et al., 2007; Bianchi et al., 2009; Maiolino et al., 2010; Sanfrutos et al., 2013; Miniutti et al., 2014), to weeks or months (e.g. Lamer et al., 2003; Marinucci et al., 2013; Jana et al., 2022; Torres-Albà et al., 2025) and to more than a year (e.g. Markowitz et al., 2014; Hernández-García et al., 2015; Laha et al., 2020; Torres-Albà et al., 2023; Pizzetti et al., 2025). Furthermore, both Seyfert 1 and Seyfert 2 galaxies display variability (e.g. Malizia et al., 1997). By studying the variability of in these objects, constraints may be placed on the obscuring material properties, such as size and location (e.g. Risaliti et al., 2002, 2005; Pizzetti et al., 2022; Marchesi et al., 2022; Torres-Albà et al., 2025). Some studies suggest that the obscuration variability may result from other geometries, such as a warped disk in the case of NGC 4258 (e.g. Fruscione et al., 2005; Masini et al., 2022), or the presence of outflows (e.g. Kaastra et al., 2014; Mehdipour et al., 2017; Kara et al., 2021; Lian et al., 2025).
However, some studies fail to find variability in a significant fraction of AGN (e.g. Hernández-García et al., 2015; Laha et al., 2020; Torres-Albà et al., 2023). The sample used in Torres-Albà et al. (2023) was selected based on data from Zhao et al. (2021) that already indicated variability (flux or ) was necessary for a good fit. Despite this, only 5 out of 12 sources were confidently classified as variable in . It is unclear how likely it is to observe a constant across many different timescales if the obscuring material is in fact clumpy and dynamic as described above. One potential explanation is that the column consists of many clouds, such that it is unlikely that the number of clouds currently in our line of sight will change significantly enough to be registered as a change in . However, Nenkova et al. (2008), for example, argues that to reproduce the observed 1-100 m SED, the average number of clouds along equatorial lines of sight can be no larger than 10–15 with individual column densities cm-2. Another explanation is that, in some galaxies, the primary source of obscuration is the host galaxy itself, rather than the compact torus (e.g. Laha et al., 2020). Some studies have indicated that Seyfert 2 galaxies are more likely to be found in galaxies with large-scale dusty structures crossing the line of sight, which may be able to explain the absorption (e.g. Malkan et al., 1998). For example, Prieto et al. (2014) looked at seven type 2 galaxies and observed in all cases, kpc-scale dust filaments crossing the nucleus that could potentially explain the nuclear obscuration without the need to invoke a pc-scale torus.
@@
In this work, we perform the first systematic search for obscuration variability in the whole of the Chandra archive. As a result, we obtain the first estimate of the number of obscuration-variable AGN in the local Universe, and present a sample of X-ray sources that are likely -variable. Studying the timescales on which absorption variability in the X-ray emission occurs can help us constrain the location and distribution of the obscuring material, be it compact or on galactic scales. However, previous studies only focus on samples of at most sources (e.g. Markowitz et al., 2014; Hernández-García et al., 2015; Laha et al., 2020), selected in a variety of different ways, and at times purposefully biased toward obscuration variability (e.g. Torres-Albà et al., 2023; Pizzetti et al., 2025). As such, it is difficult to build up a large enough, unbiased sample of changes in to make significant conclusions.
@@
In our previous work (Cox et al., 2023), we developed a method based on hardness ratios that could predict whether or not two X-ray observations of a given AGN were likely to display variability, without going through the time-consuming process of detailed spectral modeling. We apply this method to all archival Chandra data for sources with and low probability for pileup, based on their observed count rate. This allows us to select, from a relatively unbiased sample, sources for follow-up that will likely be more profitable in answering questions about the obscuring material in AGN. Furthermore, our method provides a quick estimate on the lower limit of -variable sources in the local universe, a quantity that informs the physical requirements of clumpy torus models such as UXCLUMPY (Buchner et al., 2019) and XCLUMPY (Tanimoto et al., 2019).
@@
In Section 2, we describe our initial sample selection and data reduction. In Section 3, we briefly recount the variability selection method presented in Cox et al. (2023). We present our variable sample selected by the method in Section 4. In Section 5, we discuss preliminary conclusions based on our results. We provide a summary in Section 6. Throughout the paper, we use a standard cosmology, with km s-1 Mpc-1, , and . All reported errors are at the 90 % confidence level unless stated otherwise.
@@ 2 Sample Selection and Data
Our sample selection starts from the Million Quasars (Milliquas) Catalogue, v8 (Flesch, 2023), accessed through the VizieR service (Ochsenbein et al., 2000). We first make a redshift cut and only include AGN with , to avoid distant, bright Seyfert 1 galaxies from dominating our sample statistics and also because our method has only been tested on local sources. This cut results in 24,605 sources. We then select those objects with a non-empty X-ray ID, which means the source has been detected by at least one X-ray instrument. This results in 3,493 sources. Of these, we find 917 sources have archival Chandra ACIS-S observations with no grating. We consider only those sources with at least two such observations resulting in 233 AGN. Removing 5 blazars results in 228 AGN. For each observation of these 228 AGN, we keep only those with at least 50 net counts in the keV band to ensure high quality HR measurements, resulting in 120 sources still with at least two qualifying observations.
@@
The net counts in each energy band were obtained from the counts in the source and background regions. Errors were obtained on the net counts following Gehrels (1986). When necessary, the counts in each band were added (and the errors added in quadrature) to obtain the total net counts and error in the hard and soft bands. See Section 3.1 of Cox et al. (2023) for more details and the equations used.
@@ 3 Hardness Ratios and Variability
In this section, we briefly recount the hardness ratio method outlined in Cox et al. (2023) and the criteria for flagging a source as variable.
@@ 3.1 Hardness Ratios
We use two different definitions for the H and S bands. For HR1, we define the S band to be 2–4 keV and the H band to be 4–10 keV. For HR2, we define S to be 4–6 keV and H to be 6–10 keV. These two hardness ratios were chosen to be sensitive to variability in sources that are already obscured with cm-2. At obscuration levels higher than cm-2, all the primary emission below 4 keV becomes absorbed and HR1 actually decreases with increasing . The second hardness ratio, HR2, is used to break the degeneracy that occurs in sources that vary in this range of . A visual explanation of these two hardness ratios and their different trends can be found in Figure 2 of Cox et al. (2023).
We ignore data below 2 keV for the following reasons. First, it is not particularly useful in detecting changes in in obscured sources because all the emission at these energies has already been reprocessed. While, the keV band can be more useful for sources with cm-2, our HR1 is still sensitive enough to detect large changes in this regime. Second, this data can be contaminated by photons from non-AGN processes. While this would be a negligible issue assuming a lack of intrinsic flux variability, AGN are known to be variable in flux, and thus such an assumption cannot be adopted. Finally, and most importantly, due to the degrading sensitivity of the Chandra ASIS-S detector at energies below 2 keV, the hardness ratio of a non-variable source would become harder in later cycles despite no spectral variability taking place (see Cox et al. 2023, Figure 1). Since these energies likely do not provide useful information for at least half of the sample, and would further require a correction to prevent it from affecting the detection method, we choose to ignore them entirely. However, we do report the variable fractions with and without the Seyfert 1 sources in Section 4 since they may have a higher false negative percentage due to the lower obscuration levels.
More details on the selection of the hard and soft energy bands as well as equations for the net counts and uncertainties can be found in Section 3 of Cox et al. (2023).
@@ 4 Results
Our analysis flagged 43 out of the 79 sources ( %) as variable at the threshold. The error on this fraction is at 90 % confidence level calculated using the Wilson interval. Considering only Seyfert 2 galaxies, 18 out of 38 sources are variable, resulting in a fraction of %. Considering only Seyfert 1 galaxies, 25 out of 41 are variable so %. From our previous work on a sample of 72 observation pairs over 12 sources, we found that the true positive rate (the percentage of truly variable sources flagged), TPR, and the false positive rate (the percentage of non-variable sources flagged), FPR, of this method are about 85 % and 40 %, respectively (Cox et al., 2023). These 12 AGN had been analyzed using complex, self-consistent torus models which account for the prominent reflection features at high obscuration, to obtain the “true” variability state of each observation pair. Assuming these rates will hold for our current sample, we can account for these biases with the following equation
@@
The fraction of flagged variable sources as a function of the minimum number of total observations is shown in the inset to Figure 2. The dark green line shows the fraction with the flagging threshold set to . The green and light green lines show the fractions calculated using the higher, more conservative thresholds of and , respectively. According to the results in Cox et al. (2023), the false positive rates for the higher thresholds drop to % while the true positive rates remain %. The dashed lines in Figure 2 represent the “corrected” fractions using Equation 4. The dark green dashed line () assumes TPR0.85 and FPR0.40. The green dashed line () assumes TPR0.70 and FPR0.20. The light green dashed line () assumes TPR0.60 and FPR0.05. The corrections for the higher thresholds are larger because the TPR and FPR are both lower. The assumption is that we are missing many sources (low TPR) and most of the sources we found are truly variable (low FPR). Therefore, there is a larger correction upwards, and a lower correction downwards when compared to the threshold. The corrected fractions are simply shown as a comparison to the observed fraction, based on our previous results. We are not claiming these to be absolute. Since the FPR of the most conservative threshold is very low, we claim the observed lower limit on the variable fraction to be around 50 % as this is the fraction for sources with a reasonable observational sampling (5–10 observations) for the highest threshold where the FPR should be close to zero.
@@
The 43 flagged sources are shown in Table 1 in order of increasing RA. The number of available ACIS-S, ACIS-I, XMM-Newton, and NuSTAR observations are shown with the number of observations used in this analysis in parentheses. variability was previously studied in three of these sources (NGC 833, NGC 835, and NGC 4388) by Torres-Albà et al. (2023). They classified NGC 835 and NGC 4388 as variable. NGC 833 was not classified as variable in that work due to their very strict definition of variable (i.e. requiring that no alternative spectral fits exist that provide similar fit statistics when not considering variability) although they report it does show some form of variability. variability has also previously been found in NGC 7582 (e.g., Piconcelli et al., 2007; Rivers et al., 2015; Braito et al., 2017; Laha et al., 2020). We are not aware of any studies focusing specifically on obscuration variability for any of the other sources in this sample.
@@ 5 Discussion
5.1 Number of observations
The variable fraction of % presented here should be considered a measure of the observed variability fraction given an incomplete temporal sampling of all the sources, not a measure of the true variable fraction. If a source is observed to be non-variable over a handful of observations, it may be that we simply did not observe the source at the right time. Therefore, this value should be considered a lower limit on the true variability fraction. This is hinted at by Figures 2 and 3. Figure 3 shows the distribution of the number of Chandra observations for the whole sample, but colored with the proportion of the sources that are variable. As can be seen, of the 34 sources with only two observations, just nine ( %) were found to display variability. On the other hand, all 11 sources with observations were flagged as variable. The median number of observations for the total sample is 3, while the median for the variable sample is over twice that for the nonvariable sample (4.5 and 2, respectively). The median for the non-variable sample is 2. This could indicate that the observed variable fraction will increase as more data is collected.
It could also be the case that known variable sources have more observations because they are targeted more frequently to analyze their variable nature; however, this does not appear to be the case here based on a search of the literature and proposal abstracts. Only two sources, ESO 253-3 (with 11 observations used) and PGC 214276 (with 12 observations used), were monitored specifically for expected variability. In the case of ESO 253-3, there is a suspected periodic tidal disruption event (e.g. ID:21708757, P.I. Anna Payne; ID:23700492, P.I. Katie Auchettl; ID:24700009, P.I. Norbert Schartel; ID:26700325, P.I. Katie Auchettl), and in the case of PGC 214276, they were hoping to observe a SMBH merger (ID:23708834, P.I. Ning Jiang). In both cases, the accretion properties would be changing rapidly, potentially resulting in a large change in photon index which may be the cause of our variability detection. Removing these two sources from the list of variable sources would result in a variable fraction of %. However, it is worth noting that we have no reason to believe there could be such extreme variability in photon index occurring for the rest of the sample. For example, in previous studies of variability for almost 30 type 2 sources, none required a variable photon index to achieve a good fit, although many required flux variability (Pizzetti et al., 2022; Torres-Albà et al., 2023; Pizzetti et al., 2025; Torres-Albà et al., 2025; Sengupta et al., 2025). To our knowledge, no other sources identified as variable in this study were targeted for follow-up observations based on known variability. Therefore the trend observed here is not significantly affected by this bias.
The case of PGC 94626 suggests that it is likely every source would be seen to vary given enough observations. We observe variability in this source, despite it being remarkably consistent throughout its observed period. Across 2,269 total observation pairs, only 48 ( %) of those pairs were flagged as variable. In other words, it is highly unlikely that this source would be flagged as variable if there were only 2 observations. Similarly, Laha et al. (2020) looked for variability in 20 sources and found 8 object/instrument combinations in which variability was observed. The number of observations in each of these groups were [3,3,4,4,4,6,9,10]. In contrast, among the 27 non-variable object/instrument pairs:

### v2

@@ Abstract
The nature of the obscuring material in active galactic nuclei (AGN) can be studied by measuring changes in the line-of-sight column density, , over time. This can be accomplished by monitoring AGN over long periods of time and at all timescales. However, this can only be done for a few selected objects as it is resource intensive. Therefore, the best option currently is to focus on population statistics based on the available archival data. In this work, we study 79 Seyfert 1 and Seyfert 2 galaxies from the Milliquas catalog to estimate a lower limit on the fraction of sources in the local universe that display spectral variability among observations. We find that 43 sources ( %), show indications of variability at 90 % confidence level. Interestingly, we also find that the variable fraction is similar for both Seyfert 1 ( %) and Seyfert 2 ( %) galaxies. The slightly higher fraction could be due to either a physical difference in the obscurers or the higher data quality in the Sy1 population. We also search for potential dependencies on the timescale between variable and non-variable observation pairs within a given source. In agreement with previous studies, we find evidence that more variability occurs on longer timescales than on shorter timescales. We present the 43 variable sources as a promising sample for future variability studies.
@@ 1 Introduction
Active galactic nuclei (AGN) shine brightly at all wavelengths due to the accretion of matter onto a supermassive black hole (SMBH). The unified model suggests the presence of a dusty, torus-shaped structure that obscures certain lines of sight to the central region, resulting in different AGN classifications depending on the orientation of the system with respect to the observer (e.g. antonucci_unified_1993; urry95).
@@
The nature of this obscuring material remains uncertain. Infrared observations, corroborated by theoretical considerations, suggest an inhomogeneous or clumpy medium (e.g. nenkova02; nenkova08; gandhi09; stalevski12; alonso16; garcia16). Many ideas have been proposed to explain the origin of the inhomogeneities and maintenance of the geometrical thickness including collisional heating (e.g. krolik_molecular_1988), IR radiation (e.g. Pier92; Krolik07), or heating by supernovae in starbursts (e.g. Fabian98; Wada02; Wada09). Winds driven by hydromagnetic forces (e.g. Konigl94; Elitzur06) or radiation pressure and feedback effects (e.g. Wada12; Schartmann14) have also been suggested. All of these mechanisms involve a very complex and dynamic environment, which is also responsible for the obscuration.
@@
X-ray observations further suggest inhomogeneity and dynamism through variability in the line-of-sight hydrogen column density, (e.g. Risaliti2002). Variability in has been observed on timescales ranging from less than a day (e.g. Elvis2004; Risaliti2005; Puccetti2007; Bianchi09; Maiolino2010; Sanfrutos13; Miniutti2014; torricelli2014search), to weeks or months (e.g. Lamer03; Marinucci13; Jana22; NTA25) and to more than a year (e.g. Markowitz2014; hernandez2015x; Laha2020; NTA23; Pizzetti2025). Furthermore, both Seyfert 1 and Seyfert 2 galaxies display variability (e.g. Malizia1997).
The source of this observed obscuration variability remains uncertain. It is not always clear if the spectral variability is truly due to changes in obscuration, as it has been suggested that changes in accretion could result in a source appearing to transition between compton-thin and compton-thick. This is because a recently turned off AGN would become reflection dominated for some time after, assuming the reflection comes from a parsec scale torus (e.g. Guainazzi98; Uttley99; Risaliti00; Guainazzi02; Matt09). However, many studies attribute these changes to clouds in orbit around the SMBH passing in and out of the line of sight (e.g. Warwick88; Risaliti2002; Risaliti2005; Risaliti07; Risaliti09; Risaliti10; Nardini11; Guainazzi16; Marinucci16; Pietrini19; Zaino20; pizzetti_multi-epoch_2022; marchesi_compton-thick_2022; Pietrini24; NTA25). On the other hand, some studies attribute the variable obscuration to disk winds or outflows (e.g. Risaliti11; Kaastra2014; Beuchert15; mehdipour2017; Kaastra18; Dehghanian19; Kriss19; Kara2021; Serafinelli21; Lian25; Lyu25; Peca25). A recent review on changing-obscuration AGN can be found in Ricci23.
However, most studies fail to find variability in a significant fraction of AGN (e.g. hernandez2015x; Laha2020; NTA23). The sample used in NTA23 was selected based on data from Zhao21 that already had a good indication for variability (either in intrinsic flux or ). Despite this, only 5 out of 12 sources were confidently classified as variable in . It is unclear how likely it is to systematically observe a constant in a large fraction of sources if the obscuring material is in fact clumpy. Potential explanations include very large numbers of clouds in the line of sight (however, see nenkova08, for arguments against this hypothesis) or large structures within the host galaxy being the primary obscurer (e.g. Malkan98; Prieto14; Laha2020).
@@
The literature lacks systematic searches for obscuration variability that can give an indication of how common this phenomenon is, which in turns limits our understanding of its origin and typical properties (e.g. timescales, frequency). Previous studies only focus on samples of at most sources (e.g. Markowitz2014; hernandez2015x; Laha2020), selected in a variety of different ways, and at times purposefully biased toward obscuration variability (e.g. NTA23; Pizzetti2025). As such, it is difficult to build up a large enough, unbiased sample of changes in to make significant conclusions.
@@
In this work, we perform the first systematic search for obscuration variability in the whole of the Chandra archive. As a result, we obtain the first estimate of the number of obscuration-variable AGN in the local Universe, and present a sample of X-ray sources that are likely -variable. Studying the timescales on which absorption variability in the X-ray emission occurs can help us constrain the location and distribution of the obscuring material, be it compact or on galactic scales. In our previous work (Cox2023), we developed a method based on hardness ratios that could predict whether or not two X-ray observations of a given AGN were likely to display variability, without going through the time-consuming process of detailed spectral modeling. We apply this method to all archival Chandra data for sources with and low probability for pileup, based on their observed count rate. This allows us to select, from a relatively unbiased sample, sources for follow-up that will likely be more profitable in answering questions about the obscuring material in AGN. Furthermore, our method provides a quick estimate on the lower limit of -variable sources in the local universe, a quantity that informs the physical requirements of clumpy torus models such as UXCLUMPY (buchner_x-ray_2019) and XCLUMPY (Tanimoto19).
@@
In Section 2, we describe our initial sample selection and data reduction. In Section 3, we briefly recount the variability selection method presented in Cox2023. We present our variable sample selected by the method in Section 4. In Section 5, we discuss preliminary conclusions based on our results. We provide a summary in Section 6. Throughout the paper, we use a standard cosmology, with km s-1 Mpc-1, , and . All reported errors are at the 90 % confidence level unless stated otherwise.
@@ 2 Sample Selection and Data
Our sample selection starts from the Million Quasars (Milliquas) Catalogue, v8 (Flesch2023), accessed through the VizieR service (vizier2000). We first make a redshift cut and only include AGN with , to avoid distant, bright Seyfert 1 galaxies from dominating our sample statistics and also because our method has only been tested on local sources. This cut results in 24,605 sources. We then select those objects with a non-empty X-ray ID, which means the source has been detected by at least one X-ray instrument. This results in 3,493 sources. Of these, we find 917 sources have archival Chandra ACIS-S observations with no grating. We consider only those sources with at least two such observations resulting in 233 AGN. Removing 5 blazars results in 228 AGN. For each observation of these 228 AGN, we keep only those with at least 50 net counts in the keV band to ensure high quality HR measurements, resulting in 120 sources still with at least two qualifying observations.
@@
The net counts in each energy band were obtained from the counts in the source and background regions. Errors were obtained on the net counts following gehrels_confidence_1986. When necessary, the counts in each band were added (and the errors added in quadrature) to obtain the total net counts and error in the hard and soft bands. See Section 3.1 of Cox2023 for more details and the equations used.
@@ 3 Hardness Ratios and Variability
In this section, we briefly recount the hardness ratio method outlined in Cox2023 and the criteria for flagging a source as variable.
@@ 3.1 Hardness Ratios
We use two different definitions for the H and S bands. For HR1, we define the S band to be 2–4 keV and the H band to be 4–10 keV. For HR2, we define S to be 4–6 keV and H to be 6–10 keV. These two hardness ratios were chosen to be sensitive to variability in sources that are already obscured with cm-2. At obscuration levels higher than cm-2, all the primary emission below 4 keV becomes absorbed and HR1 actually decreases with increasing . The second hardness ratio, HR2, is used to break the degeneracy that occurs in sources that vary in this range of . A visual explanation of these two hardness ratios and their different trends can be found in Figure 2 of Cox2023.
We ignore data below 2 keV for the following reasons. First, it is not particularly useful in detecting changes in in obscured sources because all the emission at these energies has already been reprocessed. While, the keV band can be more useful for sources with cm-2, our HR1 is still sensitive enough to detect large changes in this regime. Second, this data can be contaminated by photons from non-AGN processes. While this would be a negligible issue assuming a lack of intrinsic flux variability, AGN are known to be variable in flux, and thus such an assumption cannot be adopted. Finally, and most importantly, due to the degrading sensitivity of the Chandra ASIS-S detector at energies below 2 keV, the hardness ratio of a non-variable source would become harder in later cycles despite no spectral variability taking place (see Cox2023, Figure 1). Since these energies likely do not provide useful information for at least half of the sample, and would further require a correction to prevent it from affecting the detection method, we choose to ignore them entirely. However, we do report the variable fractions with and without the Seyfert 1 sources in Section 4 since they may have a higher false negative percentage due to the lower obscuration levels.
More details on the selection of the hard and soft energy bands as well as equations for the net counts and uncertainties can be found in Section 3 of Cox2023.
@@ 4 Results
Our analysis flagged 43 out of the 79 sources ( %) as variable at the threshold. The error on this fraction is at 90 % confidence level calculated using the Wilson interval. Considering only Seyfert 2 galaxies, 18 out of 38 sources are variable, resulting in a fraction of %. Considering only Seyfert 1 galaxies, 25 out of 41 are variable so %. From our previous work on a sample of 72 observation pairs over 12 sources, we found that the true positive rate (the percentage of truly variable sources flagged), TPR, and the false positive rate (the percentage of non-variable sources flagged), FPR, of this method are about 85 % and 40 %, respectively (Cox2023). These 12 AGN had been analyzed using complex, self-consistent torus models which account for the prominent reflection features at high obscuration, to obtain the “true” variability state of each observation pair. Assuming these rates will hold for our current sample, we can account for these biases with the following equation
@@
The fraction of flagged variable sources as a function of the minimum number of total observations is shown in the inset to Figure 2. The dark green line shows the fraction with the flagging threshold set to . The green and light green lines show the fractions calculated using the higher, more conservative thresholds of and , respectively. According to the results in Cox2023, the false positive rates for the higher thresholds drop to % while the true positive rates remain %. The dashed lines in Figure 2 represent the “corrected” fractions using Equation 4. The dark green dashed line () assumes TPR0.85 and FPR0.40. The green dashed line () assumes TPR0.70 and FPR0.20. The light green dashed line () assumes TPR0.60 and FPR0.05. The corrections for the higher thresholds are larger because the TPR and FPR are both lower. The assumption is that we are missing many sources (low TPR) and most of the sources we found are truly variable (low FPR). Therefore, there is a larger correction upwards, and a lower correction downwards when compared to the threshold. The corrected fractions are simply shown as a comparison to the observed fraction, based on our previous results. We are not claiming these to be absolute. Since the FPR of the most conservative threshold is very low, we claim the observed lower limit on the variable fraction to be around 50 % as this is the fraction for sources with a reasonable observational sampling (5–10 observations) for the highest threshold where the FPR should be close to zero.
@@
The 43 flagged sources are shown in Table 1 in order of increasing RA. The number of available ACIS-S, ACIS-I, XMM-Newton, and NuSTAR observations are shown with the number of observations used in this analysis in parentheses. variability was previously studied in three of these sources (NGC 833, NGC 835, and NGC 4388) by NTA23. They classified NGC 835 and NGC 4388 as variable. NGC 833 was not classified as variable in that work due to their very strict definition of variable (i.e. requiring that no alternative spectral fits exist that provide similar fit statistics when not considering variability) although they report it does show some form of variability. variability has also previously been found in NGC 7582 (e.g., Piconcelli2007; Rivers2015; Braito2017; Laha2020). We are not aware of any studies focusing specifically on obscuration variability for any of the other sources in this sample.
@@ 5 Discussion
5.1 Impact of the number of observations on the observed variable fraction
The variable fraction of % presented here should be considered a measure of the observed variability fraction given an incomplete temporal sampling of all the sources, not a measure of the true variable fraction. If a source is observed to be non-variable over a handful of observations, it may be that we simply did not observe the source at the right time. Therefore, this value should be considered a lower limit on the true variability fraction.
Figure 3 shows the distribution of the number of Chandra observations for the whole sample, but colored with the proportion of the sources that are variable. As can be seen, of the 34 sources with only two observations, just nine ( %) were found to display variability. On the other hand, all 11 sources with observations were flagged as variable. The median number of observations for the total sample is 3, but when split into the variable and non-variable samples, it becomes 4.5 and 2, respectively. This suggests that the observed variable fraction would increase if more data was available.
The case of PGC 94626 (see Section 5.6) further suggests that it is likely every source would be seen to vary given enough observations. Across 2,269 total observation pairs, only 48 ( %) of those pairs were flagged as variable. In other words, it is highly unlikely that this source would be flagged as variable if there were only 2 observations.
Similarly, Laha2020 looked for variability in 20 sources and found 8 object/instrument combinations444Variability was only considered between observations with the same instrument. There were 38 total object/instrument combinations to consider. in which variability was observed. The number of observations in each of these groups were [3,3,4,4,4,6,9,10]. In contrast, among the 30 non-variable object/instrument pairs: